### v1

# Odd fake $\mathbb{Q}$-homology quadrics exist

Abstract.
We show the existence of odd fake $\mathbb{Q}$-homology quadrics, namely of minimal surfaces $S$ of general type which have the same $\mathbb{Q}$-homology as a smooth quadric $Q\cong\mathbb{P}^{1}(\mathbb{C})^{2}$, but have an odd intersection form on $H^{2}(S,\mathbb{Z})/\mathrm{Tors}(S)$, where $\mathrm{Tors}(S)$ is the Torsion subgroup.
Our examples are provided by a special 1-dimensional family of surfaces isogenous to a product of unmixed type.
AMS Classification: 14J15; 14J25; 14J29; 14J80.
Key words: Surfaces of general type, homology quadric, fake quadric, group actions, surfaces isogenous to a product.
1. Introduction
Minimal smooth complex algebraic surfaces of general type which have the same $\mathbb{Q}$-homology algebra as a quadric $Q\cong\mathbb{P}^{1}(\mathbb{C})^{2}$ are called fake $\mathbb{Q}$-homology quadrics. They are exactly the smooth minimal complex surfaces with $q=p_{g}=0$, with second Betti number $b_{2}(S)=2$ (equivalently, with $K^{2}_{S}=8$) and with nef canonical divisor.
Their Severi lattice $Num(S)$, of divisor classes modulo numerical equivalence, has rank 2 and coincides with $H^{2}(S,\mathbb{Z})/\mathrm{Tors}(S)$; by Poincaré duality, it has a unimodular intersection form. The intersection form can be even (then the lattice is called a hyperbolic plane), or it is odd and then the intersection form is diagonalizable (see [15]).
In the paper [4], among other things, many examples were produced of such fake $\mathbb{Q}$-homology quadrics with even intersection form, but we were not able to show the existence of such surfaces with odd intersection form.
This article provides such a family of such surfaces, thus answering problems 1 and 3 raised in [4].
The family consists of surfaces $S$ isogenous to a higher product of unmixed type, which means that
$S\cong(C_{1}\times C_{2})/G,$
where $C_{i}$ is a smooth curve with $g(C_{i})\geq 2$ and $G$ is a finite group acting faithfully on $C_{1}$ and on $C_{2}$, such that the diagonal action $g(x,y):=(gx,gy)$ of $G$ on $C_{1}\times C_{2}$ is free.
In particular, the universal covering of $S$ the bidisk $\mathbb{H}\times\mathbb{H}$.
Indeed, up to now, all the known examples of fake homology quadrics have universal cover equal to the bidisk $\mathbb{H}\times\mathbb{H}$, and one may even conjecture that it must be always so.
The more narrow class of topologically fake quadrics refers to an old question raised by Hirzebruch ([12] Problem 25, see also [13], pages 779-780): does there exist a surface of general type which is homeomorphic to $\mathbb{P}^{1}\times\mathbb{P}^{1}$? *** One can ask the same question for a surface homeomorphic to the blow up $\mathbb{F}_{1}$ of $\mathbb{P}^{2}$ in one point: in this case the intersection form is odd.
Here is the precise form of our main result.
Theorem 1.1.
Consider the family of surfaces isogenous to a product of unmixed type with $q=p_{g}=0$, with group $G:=\mathfrak{A}_{5}$, and where
$C_{1}$ is the $G$-covering of $\mathbb{P}^{1}$ branched on three points with Hurwitz triple (monodromy triple) $(2,4)(3,5),(2,1,3,4,5),(1,2,3,4,5),$
$C_{2}$ is the $G$-covering of $\mathbb{P}^{1}$ branched on four points with Hurwitz quadruple (monodromy quadruple) $(1,2,3),(3,4,5),(4,3,2),(2,1,5).$
Then the corresponding family of surfaces isogenous to a product of unmixed type, $S=(C_{1}\times C_{2})/G$, is a family of fake $\mathbb{Q}$-homology quadrics of odd type.
The proof hinges on the symmetries of the curve $C_{1}$, the unique curve of genus $4$ with $\mathfrak{A}_{5}$-symmetry.
An interesting question would be to determine exactly which other of the known fake $\mathbb{Q}$-homology quadrics are of odd type (see the examples of [2], [4], [8], [6], [7]).
Acknowledgement.
Heartfelt thanks to Matthias Schütt for making the crucial observation that the surfaces in the first family of the classification list of surfaces isogenous to a product with $q=p_{g}=0$ have Torsion group of odd order.
The question about the existence of fake quadrics of odd type was raised by Jianqiang Yang in an email correspondence (in December 2022).
2. Definitions and basic properties
We shall work with projective smooth surfaces defined over the field $\mathbb{C}$.
Definition 2.1 ( $\mathbb{Q}$-homology Quadric).
Let $S$ be a smooth projective surface over $\mathbb{C}$.
The surface $S$ is called a $\mathbb{Q}$-homology quadric if $H^{*}(S,\mathbb{Q})$ has Betti numbers $b_{1}(S)=b_{3}(S)=0$, $b_{2}(S)=2$ (equivalently, $q(S)=p_{g}(S)=0$).
In turn, it will be called an even homology quadric if moreover the intersection form on $\mathrm{Num}(S)$ is even, that is, with matrix
$\left(\begin{array}[]{cc}0&1\\ 1&0\end{array}\right).$
It will be called an odd homology quadric if instead the intersection form is odd (hence diagonalizable with diagonal entries $(+1,-1)$).
For a $\mathbb{Q}$-homology quadric $S$, by the Noether formula we must have
$\chi(\mathcal{O}_{S})=1,c_{2}(S)=4,K_{S}^{2}=8.$
Also the long exact sequence of cohomology groups associated to the exponential exact sequence shows that, if $b_{1}(S)=0$,
$c_{1}\colon\mathrm{Pic}(S)\rightarrow H^{2}(S,\mathbb{Z})$
is an isomorphism. We identify these two groups and denote by $\mathrm{Tors}(S)$ their torsion subgroup. Then it follows that
$\mathrm{Num}(S)\cong\mathrm{Pic}(S)/\mathrm{Tors}(S)\cong H^{2}(S,\mathbb{Z})/\mathrm{Tors}(S).$
An even homology quadric $S$ must be minimal, since it contains no $(-1)$-curves. Hence, by surface classification, for an even homology quadric
  • either $S$ is rational, and then $S\cong\mathbb{F}_{2n}$ for $n\geq 0$;
  • or $S$ is a minimal surface of general type.
In the former case $S$ is simply connected, $\mathrm{Tors}(S)=0$ and $K_{S}$ is not ample.
In the case of an odd homology quadric
  • either $S$ is rational, and then $S\cong\mathbb{F}_{2n+1}$ for $n\geq 0$;
  • or $S$ is a, non necessarily minimal, surface of general type.
Definition 2.2 (Even fake Quadric).
Let $S$ be a smooth projective surface over $\mathbb{C}$. The surface $S$ will be called The reader should be made aware that, in the literature, sometimes $\mathbb{Q}$-homology quadrics of general type are referred to as fake quadrics, without further specification whether the intersection form is even or odd, see [13], [6], [7], [9]. an even fake quadric if it is an even homology quadric and is of general type.
Definition 2.3 (Odd fake Quadric).
Let $S$ be a smooth projective surface over $\mathbb{C}$. If $S$ is an odd homology quadric which is of general type, then either $S$ is not minimal and it is a one point blow up of a fake projective plane, or we call the surface $S$ an odd fake quadric, which means:
  1. (1)
    $S$ is minimal of general type with $K_{S}^{2}=8,p_{g}(S)=0$;
  2. (2)
    the intersection form on $\mathrm{Num}(S)$ is odd.
Therefore smooth minimal surfaces of general type with $K^{2}=8$ and $p_{g}=0$ are divided into two classes: the even and the odd fake quadrics.
Definition 2.4.
Let $S$ be a smooth projective surface. The surface $S$ is said to be isogenous to a higher product [3] if
$S\cong(C_{1}\times C_{2})/G,$
where $C_{i}$ is a smooth curve with $g(C_{i})\geq 2$ and $G$ is a finite group acting faithfully and freely on $C_{1}\times C_{2}$.
If there are respective actions of $G$ on $C_{1}$ and $C_{2}$ such that $G$ acts by the diagonal action $g(x,y):=(gx,gy)$ on $C_{1}\times C_{2}$, then $S$ is called of unmixed type.
If some element of $G$ exchanges two factors, then $S$ is called of mixed type.
Surfaces isogenous to a product of unmixed type with $p_{g}=0$, have been classified in [2] (their torsion groups have been classified in [1]).
We summarize here, for the reader’s benefit, their classification, their torsion groups $Tors(S)\cong H_{1}(S,\mathbb{Z})$, and an extra bit of information, namely the parity of the intersection form when it is known: the assertions ‘even’ were proven in [4], while the assertion ‘odd’ will be the object of our main theorem.
Theorem 2.5.
Let $S=(C_{1}\times C_{2})/G$ be a surface isogenous to a product of unmixed type, with $p_{g}(S)=0$; then $G$ is one of the groups in the Table 1 and the multiplicities $T_{1},T_{2}$ of the multiple fibres for the two natural fibrations $S\rightarrow C_{i}/G$ are listed in the table. For each case in the list we have an irreducible component of the moduli space of surfaces of general type, whose dimension is denoted by $D$. The property of being an even, respectively an odd homology quadric and the first homology group of $S$ are given in the third last column, respectively in the last column. The group identity $Id(G)$ consists of the group cardinality $|G|$ followed by the Atlas list number.
$G$ $Id(G)$ $T_{1}$ $T_{2}$ $parity$ $D$ $H_{1}(S,\mathbb{Z})$
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[2,5,5]$ $[3,3,3,3]$ odd 1 $(\mathbb{Z}_{3})^{2}\times(\mathbb{Z}_{15})$
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[5,5,5]$ $[2,2,2,3]$ 1 $(\mathbb{Z}_{10})^{2}$
$\mathfrak{A}_{5}$ $\langle 60,5\rangle$ $[3,3,5]$ $[2,2,2,2,2]$ 2 $(\mathbb{Z}_{2})^{3}\times\mathbb{Z}_{6}$
$\mathfrak{S}_{4}\times\mathbb{Z}_{2}$ $\langle 48,48\rangle$ $[2,4,6]$ $[2,2,2,2,2,2]$ 3 $(\mathbb{Z}_{2})^{4}\times\mathbb{Z}_{4}$
G(32) $\langle 32,27\rangle$ $[2,2,4,4]$ $[2,2,2,4]$ 2 $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{4}\times\mathbb{Z}_{8}$
$(\mathbb{Z}_{5})^{2}$ $\langle 25,2\rangle$ $[5,5,5]$ $[5,5,5]$ even 0 $(\mathbb{Z}_{5})^{3}$
$\mathfrak{S}_{4}$ $\langle 24,12\rangle$ $[3,4,4]$ $[2,2,2,2,2,2]$ even 3 $(\mathbb{Z}_{2})^{4}\times\mathbb{Z}_{8}$
G(16) $\langle 16,3\rangle$ $[2,2,4,4]$ $[2,2,4,4]$ even 2 $(\mathbb{Z}_{2})^{2}\times\mathbb{Z}_{4}\times\mathbb{Z}_{8}$
$D_{4}\times\mathbb{Z}_{2}$ $\langle 16,11\rangle$ $[2,2,2,4]$ $[2,2,2,2,2,2]$ 4 $(\mathbb{Z}_{2})^{3}\times(\mathbb{Z}_{4})^{2}$
$(\mathbb{Z}_{2})^{4}$ $\langle 16,14\rangle$ $[2,2,2,2,2]$ $[2,2,2,2,2]$ even 4 $(\mathbb{Z}_{4})^{4}$
$(\mathbb{Z}_{3})^{2}$ $\langle 9,2\rangle$ $[3,3,3,3]$ $[3,3,3,3]$ even 2 $(\mathbb{Z}_{3})^{5}$
$(\mathbb{Z}_{2})^{3}$ $\langle 8,5\rangle$ $[2,2,2,2,2]$ $[2,2,2,2,2,2]$ 5 $(\mathbb{Z}_{2})^{4}\times(\mathbb{Z}_{4})^{2}$
Table 1.
3. Proof of the main Theorem 1.1
Proof.
In view of the cited results in [2] and [1], it suffices to prove the assertion about the intersection form being odd.
As observed in Lemma 2.6 of [4], since for each divisor $D$
$D^{2}+D\cdot K_{S}=2p(D)-2,$
and $H^{2}(S,\mathbb{Z})/\mathrm{Tors}(S)=Num(S)$, the intersection form is even if and only if $D\cdot K_{S}$ is even, for all $D$. By unimodularity of the intersection form, the intersection form is even if and only if the class of $K_{S}$ in $Num(S)$ is divisible by 2.
In this special case, since the Torsion subgroup (see the table in Theorem 2.5) is of odd order, then the the class of $K_{S}$ is divisible by 2 if and only if $K_{S}$ is divisible by 2 (actually the same holds for each divisor class).
Using the ramification formula for $S\rightarrow(C_{1}/G)\times(C_{2}/G)=\mathbb{P}^{1}\times\mathbb{P}^{1}$, we see that, setting $m_{1},\dots,m_{s}$ to be the multiplicities of the multiple fibres in the first fibration, and $n_{1},\dots,n_{s^{\prime}}$ those in the second, then
$K_{S}=(-2+\sum_{j}(1-\frac{1}{m_{j}}))F_{1}+(-2+\sum_{i}(1-\frac{1}{n_{i}}))F_{2}.$
We use now (3) of Lemma 2.4 of [4] showing that in general $F_{j}=d_{j}\Phi_{j}$, where $d_{j}$ is the least common multiple of the multiplicities of the fibration $S\rightarrow C_{j}/G$, whose general fibre is denoted $F_{j}$.
Hence, in $Num(S)$, we have, for our special family,
$K_{S}\equiv\Phi_{1}+2\Phi_{2}.$
Then $[K_{S}]\in Num(S)$ is not 2-divisible if and only if $\Phi_{1}$ is not 2-divisible.
In turn, this is equivalent to the fact that the canonical divisor $K_{C_{1}}$, which is the pull back of $\Phi_{1}$, is not linearly equivalent to $2\theta$, where $\theta$ is a $G$-linearized line bundle on $C_{1}$.
The proof is concluded via showing the following Proposition.
Proposition 3.1.
The curve $C_{1}$ does not admit any $G:=\mathfrak{A}_{5}$-linearized Thetacharacteristic (that is, a divisor $\theta$ such $2\theta\equiv K_{C_{1}}$).
$C_{1}$ is the unique genus 4 curve with automorphism group containing $\mathfrak{A}_{5}$, canonically embedded as:
$C_{1}=\{(x_{1},\dots,x_{5})\in\mathbb{P}^{4}|\sum_{i}x_{i}=0,\sum_{i}x_{i}^{2}=0,\sum_{i}x_{i}^{3}=0\}.$
Proof.
For simplicity of notation, denote here $C_{1}$ by $C$, $K:=K_{C_{1}}$, $\mathfrak{A}_{5}$ by $G$.
$C$ is a curve of genus $g=4$, since the canonical divisor has degree $6=\frac{60}{10}$.
Observe that any $G$- invariant holomorphic 1-form descends to $C/G=\mathbb{P}^{1}$, hence $H^{0}(K)^{G}=0$.
STEP I) Since the irreducible nontrivial linear representations of $G$ have respective dimensions $3,3,4,5$ (see [11], page 221), it follows that $H^{0}(K)$ is the unique irreducible representation of dimension 4, the permutation representation
$V:=\{(x_{1},\dots,x_{5})\in{\mathbb{C}}^{5}|\sum_{i}x_{i}=0\}.$
Let $\theta$ be a Thetacharacteristic on $C$. Then $\theta$ has degree $3$ and $h^{0}(\theta)=h^{1}(\theta)\leq 2$.
ASSUMPTION : the divisor class of $\theta$ is $G$-invariant, hence we have (see [14]) the Theta-group
$1\rightarrow\mathbb{C}^{*}\rightarrow Theta(\theta)\rightarrow G\rightarrow 1,$
the group of automorphisms of the line bundle $\theta$ which are lifts of automorphisms of $G$. The above exact sequence splits if and only if $\theta$ admits a $G$-linearization, in which case $H^{0}(\theta)$ is a linear $G$-representation.
STEP II) $C$ is not hyperelliptic, since otherwise, if ${\mathcal{H}}$ is the hyperelliptic line bundle, $S^{2}(H^{0}({\mathcal{H}}))\subset H^{0}(K)$ would be a $G$- invariant subspace of dimension 3, contradicting the irreducibility of $H^{0}(K)$.
STEP III) If $\theta$ admits a $G$-linearization, then $dim(H^{0}(\theta))=0$, since otherwise again $S^{2}(H^{0}(\theta))\subset H^{0}(K)$ would be a $G$- invariant subspace of dimension $\leq 3$.
STEP IV) If $C$ has genus 4 and $\mathfrak{A}_{5}$-symmetry, then
$C=\{(x_{1},\dots,x_{5})\in\mathbb{P}^{4}|\sum_{i}x_{i}=0,\sum_{i}x_{i}^{2}=0,\sum_{i}x_{i}^{3}=0\},$
and $Aut(C)=\mathfrak{S}_{5}$.
Proof of Step IV).
The Hurwitz formula easily shows that $C/\mathfrak{A}_{5}\cong\mathbb{P}^{1}$, since the maximal order of an element of $G$ is 5.
Since $C$ is non hyperelliptic, $C$ is the complete intersection of a quadric $Q$ and of a cubic $M$ inside $\mathbb{P}^{3}=\{(x_{1},\dots,x_{5})\in\mathbb{P}^{4}|\sum_{i}x_{i}=0\}.$
Since $Q$ is unique, up to scalars, and also $M$ is unique modulo $Q$, while $G$ has no other irreducible representations of dimension 1 than the trivial one, $Q$ and $M$ must be symmetric polynomials.
Indeed, letting $\chi_{2}:=\chi\circ(g\mapsto g^{2})$, $S^{2}(H^{0}(K))=S^{2}(V)$ decomposes, having character $\frac{1}{2}(\chi^{2}+\chi_{2})$ which takes values $(10,1,2,0,0)$ on the conjugacy classes, as the sum $\mathbb{C}\oplus V\oplus U$, where $U$ is the irreducible 5-dimensional representation.
This reproves that there is only one invariant quadric, which may be chosen to be $\sum_{i}x_{i}^{2}=0$, by the Newton formulae relating Newton and elementary symmetric polynomials. The same argument shows that $M$ can be chosen as $\sum_{i}x_{i}^{3}=0$. In particular $C$ admits an $\mathfrak{S}_{5}$-symmetry, hence $Aut(C)=\mathfrak{S}_{5}$, because of the Hurwitz-Klein bound for $g=4$, $Aut(C)\leq 168$.
STEP V) Geometry of the quadric containing the canonical curve and projective representations of $G$.
Here $C\subset Q=\mathbb{P}^{1}\times\mathbb{P}^{1}$, and the last product decomposition amounts to seeing $K=L_{1}\otimes L_{2}$, where the two line bundles $L_{j}$ are corresponding to the two coordinate projections of $\mathbb{P}^{1}\times\mathbb{P}^{1}$, that is, $H^{0}(L_{1})$, respectively $H^{0}(L_{2})$ yield the two morphisms $C\rightarrow\mathbb{P}^{1}$.
Since $G$ has no subgroup of index 2, then each line bundle class $L_{j}$ is $G$-invariant, and $H^{0}(L_{j})$ is a projective representation of $G$.
Since the group of Schur multipliers $H^{2}(G,\mathbb{C}^{*})=H^{3}(G,\mathbb{Z})=H_{2}(G,\mathbb{Z})=\mathbb{Z}/2$, there is a unique central extension of $G$, equal to
$1\rightarrow\mathbb{Z}/2\rightarrow SL(2,5)\rightarrow\mathbb{P}SL(2,5)=\mathfrak{A}_{5}\rightarrow 1,$
and the projective representations of $G$ correspond to the linear representations of $SL(2,5)$.
The irreducible nontrivial linear representations of $SL(2,5)$ have respective dimensions $2,2,3,3,4,4,5,6$ (see [5], Section 38) and the centre acts non trivially on the two irreducible representations of dimension 2.
This is why $H^{0}(L_{1})\otimes H^{0}(L_{2})=V$ is indeed representation of $G$, and, since it is irreducible, $H^{0}(L_{1}),H^{0}(L_{2})$ are non isomorphic representations of $SL(2,5)$.
The above discussion shows in particular that the $G$-invariant divisor classes $L_{1},L_{2}$ do not admit a linearization.
STEP VI) Let $\theta$ be a $G$-invariant Thetacharacteristic: then $\theta\otimes L_{j}$ is also $G$-invariant, for $j=1,2$.
Now, $\theta\otimes L_{j}$ has degree $6$ but is not isomorphic to $K$, hence $H^{0}(\theta\otimes L_{j})$ has dimension equal to 3.
STEP VII) $H^{0}(\theta\otimes L_{j})$ is a 3-dimensional projective representation of $G$, hence a 3-dimensional linear representation of $SL(2,5)$; but, by the character table, the two irreducible linear representations of $G$ of dimension 3 lift to the two irreducible linear representations of $SL(2,5)$ of dimension 3.
Hence either $H^{0}(\theta\otimes L_{j})$ is a 3-dimensional irreducible linear representation of $G$, or it contains a nonzero invariant section $\sigma_{j}$.
STEP VIII) There cannot be two such invariant sections $\sigma_{1},\sigma_{2}$, because their product would yield an invariant quadric $q\in H^{0}(2K)$. We have in fact shown that there is only one invariant quadric, and this is $Q$: we reach a contradiction since $Q|C\equiv 0,$ while $q$ vanishes only on a finite set of points of $C$.
STEP IX) Hence there is a $j$ such that $H^{0}(\theta\otimes L_{j})$ is a 3-dimensional irreducible linear representation of $G$, and then we claim that $\theta\otimes L_{j}$ admits a $G$-linearization.
In fact, otherwise the associated Thetagroup equals $SL(2,5)$. $SL(2,5)$ acts on the total space of the line bundle $\theta\otimes L_{j}$, ad it suffices to show that the centre $\mathbb{Z}/2$ acts trivially.
We know that $\mathbb{Z}/2$ acts trivially on $H^{0}(\theta\otimes L_{j})$, hence also on the image of $H^{0}(\theta\otimes L_{j})\otimes\mathcal{O}_{C}\rightarrow\theta\otimes L_{j}$. As this image is Zariski dense, it follows that $\mathbb{Z}/2$ acts trivially on $\theta\otimes L_{j}$.
STEP X) Take $j$ such that $\theta\otimes L_{j}$ admits a $G$-linearization. Then, if $\theta$ admits a linearization, then also $L_{j}$ admits a linearization, contradicting Step V.
3.1. Final remark
There remains the question whether there is a $G$- invariant Thetacharacteristic $\theta$ inside $Pic^{3}(C)$. We have seen that it must be $H^{0}(\theta)=0$, and that $\theta\otimes L_{j}$ are $G$ invariant, for $j=1,2$.
Then $H^{0}(\theta\otimes L_{j})$ is either an irreducible $G$- representation of $G$, or it splits as an irreducible 2-dimensional representation of $SL(2,5)$ plus a trivial 1-dimensional summand.
Since $S^{2}(H^{0}(K))=\mathbb{C}\oplus V\oplus U$, the second alternative is not possible.
If the two representations were the same, their tensor product would be contained in $S^{2}(H^{0}(K))$ and would contain the second wedge product, which has dimension 3, a contradiction.
A computation with characters shows that the tensor product of the two distinct irreducible 3-dimensional representations is indeed isomorphic to $V\oplus U$ which is contained in $S^{2}(H^{0}(K))$, hence we do not get any contradiction in this way.